OpenStudy (anonymous):

Solve for y 1/y^3 + b/y^2 - 2b^2/y = 0

OpenStudy (anonymous):

\[1/y^3 + b/y^2 - 2b^2/y = 0\]

OpenStudy (anonymous):

\[\frac{1}{y^3}+\frac{b}{y^2}-\frac{2 b^2}{y}=0\]\[\frac{-2 b^2 y^2+b y+1}{y^3}=0 \]\[\frac{-(b y-1) (2 b y+1)}{y^3}=0 \]\[\left\{y=\frac{1}{b},y=-\frac{1}{2b}\right\} \]